How much water must be added to 20.0ml of a 60.0% saline solution in order to make a 45.0% saline solution?
Answer by stanbon(75887)

How much water must be added to 20.0 ml of a
60.0% saline solution in order
to make a 45.0% saline solution?
---------------------------------------
Equation:
salt + salt = salt
0*x + 0.60*20 = 0.45(x+20)
Multiply thru by 100 to get:
60*20 = 45x + 45*20
45x = 15*20
x = (1/3)20
x = 6 2/3 ml (amount of water needed for the mixture)
=========================
Cheers,
Stan H.